OpenStudy (anonymous):

PLEASE EXPLAIN THOROUGHLY! BEST EXPLAINED WITH CORRECT ANSWER GETS MEDAL! What is the slope of a line parallel to the line x = 5? Please explain how you arrived at your answer. ALSO Meghan has $1.10 in dimes, nickels, and pennies. she has three times as many nickels as pennies, and the number of dimes is two less than the number of pennies. How many dimes, nickels, and pennies does she have?

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

@Somy

OpenStudy (fortytherapper):

I would start out by setting up some equations Let D = Dimes, N = Nickels, and P = Pennies. If all the coins equal 1.10, what would that equation look like?

OpenStudy (anonymous):

This is my 5th grade little brother's math homework and I just completely think this is Algebra! But shouldn't the setup look a little something like this... D x N x P= $1.10

OpenStudy (anonymous):

@FortyTheRapper ^

jhonyy9 (jhonyy9):

d+n+p=1,10 $ 3n=p d=p-2 can you solve this system for d,n and p ?

OpenStudy (anonymous):

@jhonyy9 yes I understand how you answer the problem. Can you just put it in simple words so I can explain it to my little brother?

jhonyy9 (jhonyy9):

ok i will try it suppose there are not money - dollars - there are apples - red,green and yellow aples - ok. ? the total cantitite of aples cost 1,10 É d being red aples n --- green aples p --- yellow aples - not is right with apples - sorry

OpenStudy (anonymous):

You lost me. @jhonyy9

jhonyy9 (jhonyy9):

i think will be more easy understandably than you try realizing this exercise in reality phisicaly with realy pennies,dimes and nickels - i think so will be the best understandably sure

OpenStudy (anonymous):

I think so too. So can you explain it to me/my brother.

OpenStudy (anonymous):

@jhonyy9 ^

jhonyy9 (jhonyy9):

are you from USA ?

OpenStudy (anonymous):

Yes @jhonyy9.

jhonyy9 (jhonyy9):

do you have there with you these cantitie of pennies,dimes and nickels ?

OpenStudy (anonymous):

Cantitie? And emailing my little brother, so I can't show him a visual, I have to use words @jhonyy9.

jhonyy9 (jhonyy9):

d+n+p=1,10 $ 3n=p d=p-2 ok. so first you need to solve this sytem for d and p and n - do you can solve it ?

jhonyy9 (jhonyy9):

system - sorry

OpenStudy (anonymous):

I just need to explain to him how to do it. Full English.

jhonyy9 (jhonyy9):

sorry for my english - hope you willé can understanding it right so i think first you need understanding it how you can solve it - right ?

OpenStudy (anonymous):

Yes, I just would like the full thing explained to me so I can explain to my 11-year old brother.

jhonyy9 (jhonyy9):

first you need to know that these 1,10$ are formed from 3 tips of coins,money medals - ok. ?

OpenStudy (anonymous):

Yes.

jhonyy9 (jhonyy9):

dimes nickels and pennies d n p

OpenStudy (anonymous):

Yes.

jhonyy9 (jhonyy9):

so these amount of money is formed from dimes ,nickels and pennies but you know that from nickels there are 3 time more than pennies this mean that 3n=p and the number of dimes is 2 less than the number of pennies so this mean that the number of dimes so d is equal with number of pennies minus 2 yes ?

OpenStudy (anonymous):

Yes. By the way a penny i worth 1 cent. A dime, 10 cents. A nickel, 5 cents. If you didn't know that... Well... I'm sorry for you XD

jhonyy9 (jhonyy9):

thank you reality that i havent from what i can know it because i live in Europe and i not was never - again - in your contry - sorry

jhonyy9 (jhonyy9):

so do you understand it now ?

jhonyy9 (jhonyy9):

this exercise

OpenStudy (anonymous):

No, you clearly restated my question and not necessarily explained it to me. Sorry.

jhonyy9 (jhonyy9):

ok,. than how will be with your little brither ?

OpenStudy (anonymous):

If I don't understand, he won't understand.

jhonyy9 (jhonyy9):

brother - sorry

jhonyy9 (jhonyy9):

but you dont understand it again too ?

OpenStudy (anonymous):

ANSWER THIS PLEASE!

OpenStudy (anonymous):

What is the slope of a line parallel to the line x = 5? Please explain how you arrived at your answer.

OpenStudy (anonymous):

@fluttershyK, are you my savior?

OpenStudy (anonymous):

Do the first one please!!!

OpenStudy (fluttershyk):

P= number of pennies n= number of nickels d= number of dines we have: (1) 10d+5n+p=110 (2) n=3p (3) d= p-2 substitute (2) and (3) in (1) 10(p-2)+5(3p)+p=10 or: 10p-20+15p+p=110 26p=130 p=5 from (2); n=15 from (3); d=3

OpenStudy (anonymous):

Now the first question?

OpenStudy (anonymous):

@fluttershyK ^

OpenStudy (fluttershyk):

@JacobLHoltz yes what it is it

OpenStudy (anonymous):

What is the slope of a line parallel to the line x = 5? Please explain how you arrived at your answer.

OpenStudy (anonymous):

@fluttershyK ^

OpenStudy (fluttershyk):

ok I'm not sure on that one

OpenStudy (anonymous):

Can you try and help @fluttershyK?

OpenStudy (fluttershyk):

I have no idea maybe you need to go back and read your lesson or ask your teacher I'm sorry I only know so much I wish I could help more but that one I have no idea

OpenStudy (anonymous):

@AloneS who do you think can help?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

honestly the answers simple 15 nickels 3 dimes and 5 pennies.